# Narrowing the Focus: Learned Optimizers for Pretrained Models

### v2

@@ 1 Introduction
Learned optimizers or learning-to-learn methods (Schmidhuber, 1987; Bengio et al., 1990, 2013) offer a potential solution to this challenge by automating the optimizer selection process. These methods automate this process by either learning the hyperparameters of existing optimizers or developing entirely new optimization algorithms. However, two significant challenges limit widespread adoption of learned optimizers: meta-optimization difficulties, and meta-generalization.
@@
Meta-generalization refers to a learned optimizer’s ability to perform well on novel tasks it was not specifically trained for. One appraoch to achieving this is to train the optimizer on the widest possible range of tasks. Previous research focused on developing a single, adaptable algorithm by training it on a very broad domain. For example, Versatile Learned Optimizers (VeLO, Metz et al., 2022) was trained on hundreds of different problems, ranging from simple linear regression to reinforcement learning. However, this “one-size-fits-all” approach has significant drawbacks. The computational cost of training such an optimizer is immense, requiring 4 00040004\,0004 000 TPU-months in the case of VeLO. Furthermore, a single optimizer struggles to effectively adapt to the vast array of loss surfaces encountered across such a diverse set of tasks. This makes it challenging to achieve consistently high performance across all tasks.
Instead of aiming for broad applicability, we focus on the increasingly relevant domain of fine-tuning pretrained models. This specialization allows us to develop learned optimizers in a more targeted way. By narrowing the scope of the meta-training domain, the learned optimizer can specialize and excel in the specific types of tasks it is designed for. Fine-tuning also simplifies the optimization problem by leveraging existing knowledge encoded in pretrained checkpoints (Wei et al., 2021) and is typically done for a relatively few iteration, since only a small alignment with the current task is needed to achieve a good results.
To achieve this, we introduce L3RS (Learned Layer-wise Learning Rate Scheduler, pronounced “lers”), a novel learned optimizer designed specifically to leverage the performance of a set of predefined base optimizers within a narrower domain. L3RS uses a variety of performance features, such as adaptive exponential moving averages (EMA) across different scales. EMAs, or model averaging, have demonstrated improved generalization in real-world applications (Tarvainen & Valpola, 2017; Izmailov et al., 2018). Based on these features, L3RS produces layer-wise updates as a linear combination of the predefined base optimizers, along with a corresponding learning rate for each layer.
@@
  • We demonstrate that meta-training a learned optimizer on a narrow domain results in an optimizer that can outperform existing methods, with more that 50% speedup compared to another learned optimizer baseline, and more than 2×2$\times$2 × speedup compared to best performing traditional optimizers.
  • We evaluate the meta-generalization of the proposed optimizer across several components: generalization to longer training horizon, different pretraining and model initializations, and different evaluation dataset.
@@ 2 Preliminaries
Figure 1: Left: Inner loop evaluation. Given a task $T:=\left\{\theta_{0},\{D^{t}_{K}\},D^{e}\right\}$, the learned optimizer $f_{\psi}$ uses optimization statistics $\Phi$ to update model parameters starting from $\theta_{0}$ for each of $\{D^{t}_{K}\}$ batches. After $K$ update steps, the final model parameters are evaluated on the evaluation set $D^{e}$ using meta-loss $L_{\psi}$. Right: Outer loop NES meta-training iteration. Given a task $T$ and meta-parameters $\psi_{T}$, Gaussian noise is added to $\psi_{t}$ to produce a number of candidates equal to the population size $c$, $(\psi_{t,0},...,\psi_{t,c})$. An inner loop evaluation is performed on the given task for all candidates. The fitness of each candidate is then used to perform an NES update step, resulting in the next learned optimizer parameters, $\psi_{t+1}$.
Consider a loss function $L(\theta,X)$, parameterized by $\theta$ for a given data $X$. Learned optimizers $f_{\psi}$, parameterized by $\psi$, provide a direction to update the model’s weights $\theta$ as
@@
where $\Phi$ is a collection of features that describes the current (or past) optimization statistics, such as the gradient or the loss value. For example, the simplest learned optimizer might simply optimize the learning rate $\lambda$ of SGD, as $f_{\psi}(\Phi;\theta_{n},X)=\lambda\nabla L(\theta_{n},X)$, in which case $\Phi:=\{\nabla L\}$ and $\psi:=\{\lambda\}$. In case of Adam, the optimization parameters would be $\psi:=\{\lambda,\beta_{1},\beta_{2}\}$.
@@
Learned optimizer parameters $\psi$ are evaluated by performing $K$ update steps (1) on every batch from $D^{t}_{K}$, to get $\theta_{K}$ and calculating the loss using the evaluation batch $L_{\psi}(\theta_{K},D^{e})$.
@@
After the learned optimizer is trained, it can be evaluated to a new task distribution $\widetilde{T}:=\{\widetilde{\theta}_{0},\{\widetilde{D}^{t}_{K}\},\widetilde{D}^{e}\}$.
@@ 3 Motivation and Related Work
We can separate the optimizers into two main categories. First category, so called black-box optimizers (Li & Malik, 2016; Andrychowicz et al., 2016; Wichrowska et al., 2017; Lv et al., 2017; Sandler et al., 2021; Metz et al., 2020a, b, 2022), learns an update function $f_{\psi}$ from scratch using a custom inner optimization loop. These methods often have a large number of parameters, making them prone to overfitting and stability issues (Harrison et al., 2022). While recent work has explored using Transformers as a meta-learner architecture (Chen et al., 2022; Moudgil et al., 2023; Jain et al., 2024), they have not yet shown a significant advantage over traditional optimizers or other learned optimizer architectures.
A notable example of black-box optimizer is VeLO, which uses per-tensor HyperNetworks (Ha et al., 2016) represented by 512-wide LSTMs to generate parameters of per-parameter MLPs. To compute the parameter updates, it passes the per-parameter features through the generated MLP network consisting of 2-hidden layer, 4-hidden MLP. The input to the HyperNetwork and MLP vary slightly, but generally represent the training dynamic of the optimization. VeLO is designed to be general and has a lot of parameters that need to be learned, which makes meta-training very expensive. We envision an ideal optimizer would be lean and ideally fast to train on a narrow task domain.
@@
The second category of meta-optimizers learns a higher-level meta-component on top of existing hand-designed optimizers. Such approaches include learning to adapt the hyperparameters of an underlying optimizer Daniel et al. (2016), or learn a schedule for the learning rate (Xu et al., 2017, 2019). In addition, these methods can include adapting the optimizer’s parameters directly (Shaban et al., 2019). Hypergradient methods (Maclaurin et al., 2015; Baydin et al., 2017; Grazzi et al., 2020; Moskovitz et al., 2019) perform a hyperparameter search on-the-fly during optimization.
Our proposed approach strikes a balance between flexibility of black-box minimizer and the stability of learning hyperparameters of known optimizers. We leverage the benefits of hand-designed optimizers while incorporating the adaptability of learned components using a simple small neural network for learning the per-layer learning rate. Unlike existing methods that learn a fixed set of hyperparameters for a single existing optimizer, L3RS learns from a set of base-optimizers. Moreover, during meta-training we also search over the hyperparameters of the base-optimizer.
@@
Similar to our paper Landro et al. (2021) also propose an algorithm that combines Adam and SGD as a linear combination of their update direction. However, their method differs significantly in two key aspects: (1) they do not provide a mechanism for learning the mixing coefficient, instead treating it as fixed hyperparameters, and (2) their combination is applied globally across all layers, while our proposed method, L3RS, employs a layer-wise adaptive strategy.
@@ 4 L3RS: Learned Layer-wise Learning Rate Scheduler
Figure 2: Left: L3RS applied to a single layer of the target model. The MLP receives the norm of each direction provided, time features, target layer embedding, and EMA features of the global loss, layer gradient norm and layer parameter norm. The MLP outputs the weighting for each direction ($\mu_{p}$) as well as the final update norm ($\lambda$). Right: L3RS is applied at every layer of the network independently with the same MLP weights but different layer embeddings.
@@ 4.1 Model architecture
Figure 2 shows the main input and output components of L3RS as well is provide a flow-chart on how the weights of the underlying model $\theta$ are updated.
@@
In order to capture both short-term and log-term performance trends, for each layer $l$, we maintain four EMAs with different smoothing factors $\gamma\in\{0,0.9,0.99\}$. EMAs are updated recursively as $a^{(l)}_{i,k+1}:=\gamma_{i}a^{(l)}_{i,k}+(1-\gamma_{i})\xi^{(l)}$, where $a^{(l)}_{i,k}$ denotes the $i$th EMA for layer $l$ at time step $k$, and $\xi^{(l)}$ represents a layer-specific statistic of interest. Specifically, for a given layer we track the following:
  • Log $l_{2}$-norm of weights. $\log(\left\lVert w^{(l)}\right\rVert_{2})$, where $w^{(l)}$ are the weights for layer $l$..
@@
For each base optimizer $d_{p}$, the log $l_{2}$-norm of the direction is provided: $\log(\left\lVert d_{p}\right\rVert_{2})$.
@@ 4.2 Comparison with VeLO
Optimizer State Memory Update Function Size
SGD $\times0$ 61.4 KiB
Adam 2.0×2.0$\times$2.0 × 240.0 KiB
L3RS 2.0×2.0$\times$2.0 × 2.9 MiB
VeLO 4.0×4.0$\times$4.0 × 418.78 MiB
Table 1: Compute and memory costs for various optimizers. State memory is the ratio of the optimizer state size to the model parameter size. Update Function Size is the size of the compiled code to calculate the weight update given the gradients. Target model is ResNet-34 with 1000 classes.
@@ 5 Experiments
To evaluate the performance of our proposed L3RS optimizer, we designed a series of image classification experiments using the ResNet-34 model.
@@
During inference, the memory footprint of this L3RS optimizer state is 2×2$\times$2 × the model parameters (all from Adam). VeLO uses a memory state of 4×4$\times$4 × the model parameters (plus additional memory from $3$ AdaFactor style accumulators). Table 1 shows a comparison of the memory state requirements for various optimizers, as well as the compute cost for each.
@@
For NES training, we use a smooth exponential decay for both meta-learning rate $\alpha$ and the Gaussian noise standard deviation $\sigma$ by 0.5 every 500 generations. We use the same antithetic sampling and fitness transformations as in Salimans et al., 2017. We use a population size of $32$, meta-batch size of $4$ and train for $2000$ generations for all our experiments (except the ablation study). We parallelize the evaluation of each candidate in the population, using a total of $32$ A100 GPUs for four days.
@@ A $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
Figure 3: Meta-evaluation of the L3RS, meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks. Performance is compared to VeLO (left), Adam with cosine learning rate (center), and Adam with a constant learning rate (right). Each marker represents model evaluation at that number of steps. Solid lines indicate the number of steps for in-distribution evaluation, while dashed lines indicate generalization to more steps than meta-training.
@@
F. Speedup of L3RS. For in-domain generalization, this shows how much faster L3RS achieves a given accuracy compared to the baselines.
@@
For optimizers which parameters depend on the number of training steps $K$, such as Adam (cosine), VeLO (ft) and L3RS, we perform separate evaluations across various $K$ values to assess their performance both within and beyond the meta-training regime.
For all meta-evaluations we average the results of 100 tasks sampled from the specified task distribution. The average and standard deviation of eval accuracy across the sampled tasks is reported.
@@ Results.
Interestingly, changing either the initialization or evaluation dataset to Places has a minimal effect on the relative perfoamnce of the optimizers, even though Places was not included in the meta-training data. This highlights the robust generalization capabilities of both L3RS and the fine-tuned VeLO.
@@
In Figure 3 (F) we show the speedup, in training steps, L3RS achieves over baseline methods to reach a given accuracy. L3RS robustly is able to demonstrate 50% speedup over VeLO (ft) and $100\%$ speedup over the best hand-designed optimizer.
Finally, Figure 4 shows the learned parameters of L3RS during a $100$-step evaluation run. Interestingly, the optimizer initially heavily favors Adam direction, switching towards SGD as training progresses (with a curious sudden switch for some layers around iteration $10$). Moreover, some layers consistently rely on SGD or Adam direction throughout the entire run. While the precise interpretation of these parameter dynamics is challenging, the parameters of L3RS are considerably more interpretable than those of most black-box learned optimizers.
Base-optimizers No Embedding Per-layer MLP Embedding
SGD only $64.48\pm 5.05$ $64.71\pm 5.00$ $68.18\pm 4.29$
Adam only $61.44\pm 5.14$ $67.14\pm 4.57$ $68.52\pm 4.25$
SGD, Adam $65.70\pm 4.91$ $68.52\pm 4.40$ $\mathbf{68.93\pm 4.58}$
Table 2: Average and standard deviation evaluation accuracy given different optimizers and per-layer strategies.

### v3

@@ 1 Introduction
Learned optimizers or learning-to-learn methods (Schmidhuber, 1987; Bengio et al., 1990; 2013) offer a potential solution to this challenge by automating the optimizer selection process. These methods automate this process by either learning the hyperparameters of existing optimizers or developing entirely new optimization algorithms. However, two significant challenges limit widespread adoption of learned optimizers: meta-optimization difficulties, and meta-generalization.
@@
Meta-generalization refers to a learned optimizer’s ability to perform well on novel tasks it was not specifically trained for. One approach to achieving this is to train the optimizer on the widest possible range of tasks. Previous research focused on developing a single, adaptable algorithm by training it on a very broad domain. For example, Versatile Learned Optimizers (VeLO, Metz et al., 2022) was trained on hundreds of different problems, ranging from simple linear regression to reinforcement learning. However, this “one-size-fits-all” approach has significant drawbacks. The computational cost of training such an optimizer is immense, requiring 4 00040004\,0004 000 TPU-months in the case of VeLO. Furthermore, a single optimizer struggles to effectively adapt to the vast array of loss surfaces encountered across such a diverse set of tasks. This makes it challenging to achieve consistently high performance across all tasks.
Instead of aiming for broad applicability, we focus on the increasingly relevant domain of fine-tuning pretrained models. This specialization allows us to develop learned optimizers in a more targeted way. By narrowing the scope of the meta-training domain, the learned optimizer can specialize and excel in the specific types of tasks it is designed for. Fine-tuning also simplifies the optimization problem by leveraging existing knowledge encoded in pretrained checkpoints (Wei et al., 2021) and is typically done for a relatively few iterations, since only a small alignment with the current task is needed to achieve good results.
To achieve this, we introduce L3RS (Learned Layer-wise Learning Rate Scheduler, pronounced “lers”), a novel learned optimizer designed specifically to leverage the performance of a set of predefined base optimizers within a narrower domain. L3RS uses a variety of performance features, such as adaptive exponential moving averages (EMA) across different time scales. EMAs, or model averaging, have demonstrated improved generalization in real-world applications (Tarvainen & Valpola, 2017; Izmailov et al., 2018). Based on these features, L3RS produces layer-wise updates as a linear combination of the predefined base optimizers, along with a corresponding learning rate for each layer.
@@
  • We demonstrate that meta-training a learned optimizer on a narrow domain results in an optimizer that can outperform existing methods, with more that $50\%$ speedup compared to another learned optimizer baseline, and more than $200\%$ speedup compared to best performing traditional optimizers.
  • We evaluate the meta-generalization of the proposed optimizer across several components: generalization to longer training horizon, different pretraining and model initializations, and a different evaluation dataset.
@@ 2 Preliminaries
Figure 1: Left: Inner loop evaluation. Given a task $T:=\left\{\theta_{0},\{D^{t}_{K}\},D^{e}\right\}$, the learned optimizer $f_{\psi}$ uses optimization statistics $\Phi$ to update model parameters starting from $\theta_{0}$ for each of $\{D^{t}_{K}\}$ batches. After $K$ update steps, the final model parameters are evaluated on the evaluation set $D^{e}$ using meta-loss $L_{\psi}(\theta_{K},D^{e})$. Right: Outer loop NES meta-training iteration. Given a task $T$ and meta-parameters $\psi_{T}$, Gaussian noise is added to $\psi_{t}$ to produce a number of candidates equal to the population size $c$, $(\psi_{t,0},...,\psi_{t,c})$. An inner loop evaluation is performed on the given task for all candidates. The fitness of each candidate is then used to perform an NES update step, resulting in the next learned optimizer parameters, $\psi_{t+1}$.
Consider a loss function $L(\theta_{n},X)$, parameterized by weights $\theta_{n}$ for a given data $X$ at cetain optimization step $n$. Learned optimizers $f_{\psi}$, parameterized by $\psi$, provide a direction to update the model’s weights $\theta_{n}$ as
@@
where $\Phi$ is a collection of features that describes the current (or past) optimization statistics, such as the gradient or the loss value. For example, the simplest learned optimizer might simply optimize the learning rate $\lambda$ of SGD, as $f_{\psi}(\Phi;\theta_{n},X)=\lambda\nabla L(\theta_{n},X)$, in which case $\Phi:=\{\nabla L\}$ and $\psi:=\{\lambda\}$. In case of Adam, the optimization parameters would be $\psi:=\{\lambda,\beta_{1},\beta_{2}\}$, where $\lambda$ is the learning rate, $\beta_{1}$ and $\beta_{2}$ are the exponential decay rates for the first and second moments.
@@
Learned optimizer parameters $\psi$ are evaluated by performing $K$ update steps (equation 1) on every batch from $D^{t}_{K}$, to get $\theta_{K}$ and calculating the loss using the evaluation batch $L_{\psi}(\theta_{K},D^{e})$.
@@
After the learned optimizer is trained, it can be evaluated on a new task distribution $\widetilde{T}:=\{\widetilde{\theta}_{0},\{\widetilde{D}^{t}_{K}\},\widetilde{D}^{e}\}$.
@@ 3 Motivation and Related Work
We can separate the optimizers into two main categories. First category, so called black-box optimizers (Li & Malik, 2016; Andrychowicz et al., 2016; Wichrowska et al., 2017; Lv et al., 2017; Sandler et al., 2021; Metz et al., 2020a; b; 2022), learns an update function $f_{\psi}$ from scratch using a custom inner optimization loop. These methods often have a large number of parameters, making them prone to overfitting and stability issues (Harrison et al., 2022). While recent work has explored using Transformers as a meta-learner architecture (Chen et al., 2022; Moudgil et al., 2023; Jain et al., 2024), they have not yet shown a significant advantage over traditional optimizers or other learned optimizer architectures.
A notable example of black-box optimizer is VeLO, which uses per-tensor HyperNetworks (Ha et al., 2016) represented by 512-wide LSTMs to generate parameters of per-parameter multi-layer perceptrons (MLPs). To compute the parameter updates, it passes the per-parameter features through the generated MLP network consisting of 2-hidden layer, 4-hidden MLP. The input to the HyperNetwork and MLP vary slightly, but generally represent the training dynamic of the optimization. VeLO is designed to be general and has a lot of parameters that need to be learned, which makes meta-training very expensive. We envision an ideal optimizer would be lean and ideally fast to train on a narrow task domain.
@@
The second category of meta-optimizers learns a higher-level meta-component on top of existing hand-designed optimizers. Such approaches include learning to adapt the hyperparameters of an underlying optimizer Daniel et al. (2016), or learn a schedule for the learning rate (Xu et al., 2017; 2019). In addition, these methods can include adapting the optimizer’s parameters directly (Shaban et al., 2019). Hypergradient methods (Maclaurin et al., 2015; Baydin et al., 2017; Grazzi et al., 2020; Moskovitz et al., 2019) perform a hyperparameter search on-the-fly during optimization.
Our proposed approach strikes a balance between flexibility of black-box minimizer and the stability of learning hyperparameters of known optimizers. We leverage the benefits of hand-designed optimizers while incorporating the adaptability of learned components using a simple small neural network for learning the per-layer learning rate. Unlike existing methods that learn a fixed set of hyperparameters for a single existing optimizer, L3RS learns to utilize a set of base-optimizers. Moreover, during meta-training we also search over the hyperparameters of the base-optimizers.
@@
Similar to our paper Landro et al. (2021) also propose an algorithm that combines Adam and SGD as a linear combination of their update direction. However, their method differs significantly in two key aspects: (1) they do not provide a mechanism for learning the mixing coefficient, instead treating it as fixed hyperparameters, and (2) their combination is applied globally across all layers, while our proposed method employs a layer-wise adaptive strategy.
@@ 4 L3RS: Learned Layer-wise Learning Rate Scheduler
Figure 2: Left: L3RS applied to a single layer of the target model. The MLP receives time features, EMA features of the global loss, layer gradient norm and layer parameter norm, the target layer embedding, and the norm of each direction provided. The MLP outputs the weighting for each direction ($\mu_{p}$) as well as the final update norm ($\lambda$). Right: L3RS is applied at every layer of the network independently with the same MLP weights but different layer embeddings.
@@ 4.1 Model architecture
Figure 2 shows the main input and output components of L3RS as well as provides a flow-chart on how the weights of the underlying model $\theta$ are updated.
@@
In order to capture both short-term and log-term performance trends, for each layer $l$, we maintain three EMAs with different smoothing factors $\gamma\in\{0,0.9,0.99\}$. EMAs are updated recursively as $a^{(l)}_{i,k+1}:=\gamma_{i}a^{(l)}_{i,k}+(1-\gamma_{i})\xi^{(l)}$, where $a^{(l)}_{i,k}$ denotes the $i$th EMA for layer $l$ at time step $k$, and $\xi^{(l)}$ represents a layer-specific statistic of interest. Specifically, for a given layer we track the following:
  • Log $l_{2}$-norm of weights. $\log(\left\lVert w^{(l)}\right\rVert_{2})$, where $w^{(l)}$ are the weights for layer $l$.
@@
For each base optimizer, we provide the log $l_{2}$-norm of its update: $\log(\left\lVert d_{p}\right\rVert_{2})$.
@@ 4.2 Comparison with VeLO
Optimizer Memory Overhead Compute Overhead
SGD $\times0$ $0.003\%$
Adam 2.0×2.0$\times$2.0 × $0.019\%$
L3RS 2.0×2.0$\times$2.0 × $0.057\%$
VeLO 4.0×4.0$\times$4.0 × $0.920\%$
Table 1: Memory and compute overhead for various optimizers using a ResNet-34 model with 25 output classes. Memory Overhead is the ratio of optimizer state size to model parameter size. Compute overhead is the compute cost ratio of the optimizer update step to a full training step for a batch of 64 images. Compute cost is calculated using Jax compilation statistics.
@@ 5 Experiments
To evaluate the performance of our proposed L3RS optimizer, we designed a series of fine-tuning image classification experiments using the ResNet-34 model.
Fine-Tuning.
Our experiments focus on fine-tuning pretrained models. In particular, we are interested in optimization horizons on the order of hundreds or thousands of steps. While training can certainly extend further, this training regime is relevant in cases where the amount of training data or high computational costs can be a limiting factor.
@@
During inference, the memory footprint of this L3RS optimizer state is 2×2$\times$2 × the model parameters (all from Adam). VeLO uses a memory state of 4×4$\times$4 × the model parameters (plus additional memory from $3$ AdaFactor style accumulators). Table 1 shows a comparison of the memory state requirements for various optimizers, as well as the compute overhead for each.
@@
For NES meta-training, we use a smooth exponential decay for both meta-learning rate $\alpha$ and the Gaussian noise standard deviation $\sigma$ by 0.5 every 500 generations. We use the same antithetic sampling and fitness transformations as in Salimans et al., 2017. We use a population size of $32$, meta-batch size of $4$ and train for $2000$ generations for all our experiments (except the ablation study). We parallelize the evaluation of each candidate in the population, using a total of $32$ A100 GPUs for four days.
@@ A $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
Figure 3: Meta-evaluation of L3RS meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks. Performance is compared to VeLO (left), Adam with cosine learning rate (center), and Adam with a constant learning rate (right). Each marker represents model evaluation at that number of steps. Solid lines indicate the number of steps for in-distribution evaluation, while dashed lines indicate generalization to more steps than meta-training.
@@
F. Speedup of L3RS in iterations. For in-domain generalization, this shows how much faster L3RS achieves a given accuracy compared to the baselines.
@@
For optimizers which depend on the number of training steps $K$, such as Adam (cosine), VeLO (ft) and L3RS, we perform separate evaluations across various $K$ values to assess their performance both within and beyond the meta-training regime.
For all meta-evaluations we average the results of 100 tasks sampled from the specified task distribution. The average and standard deviation of eval accuracy across the sampled tasks is reported. All figures are reproduced using loss instead of accuracy in Appendix A.
@@ Results.
Interestingly, changing either the initialization or evaluation dataset to Places has a minimal effect on the relative performance of the optimizers, even though Places was not included in the meta-training data. This highlights the robust generalization capabilities of both L3RS and the fine-tuned VeLO.
@@
Figure 4: Visualization of learned mixing coefficients $\mu^{(l)}$ and per-layer learning rates $\lambda^{(l)}$ over 100 steps for a ResNet-34 model. Each layer’s type and component are distinguished by color and line type (see legend). The general trend shows curves moving up and to the left, indicating a transition from Adam ($\mu^{(l)}=0$) to SGD ($\mu^{(l)}=1$) and a decrease in $\lambda^{(l)}$. The initial step is marked with a $\bullet$ and the final step with a $\star$.
Figure 5: Average learning rate vs. direction mix between Adam and SGD for each of the 100 steps of the L3RS optimizer.
In Figure 3 (F) we show the speedup, in training steps, L3RS achieves over baseline methods to reach a given accuracy. L3RS is able to demonstrate robust 50% speedup over VeLO (ft) and $100\%$ speedup over the best hand-designed optimizer.
The learned parameters of L3RS exhibit interesting dynamics during a 100-step evaluation (Figure 4). Initially, the optimizer strongly favors the Adam direction with a high learning rate for most layers. As training progresses, this preference shifts, transitioning towards SGD and a lower learning rate (represented by a movement towards the top-left of the plot). Figure 5, showing the average parameter movement across all layers, reveals several distinct phases: an initial warm-up period with an increasing learning rate, a period of relatively constant learning rate while transitioning from Adam to SGD, a phase of rapid learning rate decay, and a final convergence to the SGD direction over the last $\sim$10 steps. While the precise interpretation of these parameter dynamics is challenging, the L3RS parameters are considerably more interpretable than those of most black-box learned optimizers.
Base-optimizers Embedding No Embedding Per-layer MLP Global
SGD only $68.18\pm 4.29$ $64.48\pm 5.05$ $64.71\pm 5.00$ $64.37\pm 4.94$
Adam only $68.52\pm 4.25$ $61.44\pm 5.14$ $67.14\pm 4.57$ $63.45\pm 4.84$
SGD, Adam $\mathbf{68.93\pm 4.58}$ $65.70\pm 4.91$ $68.52\pm 4.40$ $66.58\pm 4.93$
Table 2: Average and standard deviation of evaluation accuracy for different optimizers and per-layer strategies.